Solve using the Quadratic Formula:
2x^2 = 4 + 7x

Answer by nerdybill(7384)

2x^2 = 4 + 7x
2x^2 - 7x = 4
2x^2 - 7x - 4 = 0
(2x+1)(x-4) = 0
.
x = {-1/2, 4}